# 11footnotetext: co-authors22footnotetext: co-senior authors

### v1

@@ 6 Conclusion and Limitations
We demonstrate that the DGM can autonomously achieve performance on par with openly available solutions. However, it still falls short of closed-source SoTA SWE-bench solutions. An open question is whether running the DGM for longer would continue to yield performance gains and eventually surpass closed-source solutions. These closed-source solutions often rely on elaborately handcrafted techniques developed by teams of highly skilled experts. Since FMs have yet to match the capabilities of such experts (e.g., in reasoning), the DGM currently requires extensive compute to discover improvements. A single run of the DGM on SWE-bench, as presented in Section˜4, takes about 2 weeks and incurs significant API costs (Section˜B.1). We hypothesize that further progress will require more efficient use of computational resources and the development of better reasoning skills.
@@
Since this version of the DGM is mainly powered by FMs, it is inherently limited by the capabilities of the underlying FM. Hence, an exciting future direction is to extend self-modification beyond just prompts or FM workflows, to include more computationally intensive methods, such as rewriting its own training script to update the FM itself. While this version of the DGM focuses on coding, AI systems are increasingly applied across a wide range of domains (e.g., computer vision, creative writing). Another promising extension is to develop self-improving AI systems capable of enhancing themselves beyond just the coding domain. A key assumption in this work is that coding benchmarks are a good reflection of the agent’s ability to self-improve, since the self-modification task requires the agent to modify its own codebase. However, one could envision an alternative approach that co-evolves the target task distribution, thereby removing the constraint of self-improvement being tied to a single objective, as in true open-ended processes. As discussed in Section˜5, we must also continue to keep safety front and center as we explore this powerful technology.
@@
In conclusion, the DGM represents a significant step toward the automation of AI development through self-improving systems capable of editing their own codebase. While current limitations in compute and reasoning constrain its full potential, continued advances in FMs and infrastructure may unlock more powerful and general-purpose self-improvements. Provided that the safety concerns are carefully navigated (Section˜5), the future of self-improving AI systems and AI-Generating Algorithms [20] holds immense promise to open-endedly evolve AI, continually rewriting or retraining itself in pursuit of greater capabilities aligned with human values.
Acknowledgments and Disclosure of Funding
@@
This research was supported by the Vector Institute, the Canada CIFAR AI Chairs program, a grant from Schmidt Futures, an NSERC Discovery Grant, and a generous donation from Rafael Cosman. Resources used in preparing this research were provided, in part, by the Province of Ontario, the Government of Canada through CIFAR, and companies sponsoring the Vector Institute (https://vectorinstitute.ai/partnerships/current-partners/). Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the sponsors. We also thank Aaron Dharna, Ben Norman, Cédric Colas, and Shyam Sudhakaran for insightful discussions and feedback.
Supplementary Material
Appendix A Algorithmic Details
A.1 Initial Coding Agent
@@
A.2 Parent Selection
At each DGM iteration, we select a subset of agents from the archive as parents to self-modify and produce new child agents (Section˜3). The details of the parent selection process are outlined below.
@@ A.2 Parent Selection
A.3 Self-Improve Prompts
For a chosen coding agent to self-modify, we feed the evaluation logs from its performance on the coding benchmark into an FM to diagnose the next potential improvement (Section˜3). We chose to use OpenAI’s o1 model [130] for this purpose, as we found it had better reasoning capabilities than other FMs available at the time of our experiments. The prompts used for this diagnosis are provided below. From the diagnostic output, the implementation_suggestion and problem_description are extracted and passed to the coding agent as its self-modification task.
@@ A.3 Self-Improve Prompts
Self-Improve Prompt # Coding Agent Summary - **Main File**: `coding_agent.py` - Primary Class: `AgenticSystem` - The `forward()` function is the central entry point. - Prompts are located either within the `forward()` function or in the `prompts/` directory. - **Tools**: `tools/` - The `tools/` directory contains various tools that LLMs can use to perform specific tasks. {internallinenumbers*} - Each tool must have a `tool_info()` function that returns a JSON object containing ’name’, ’description’, and ’input_schema’. The ’input_schema’ should be a JSON object containing ’type’, ’properties’, and ’required’. {internallinenumbers*} - Each tool must have a `tool_function()` function that takes the arguments defined in input_schema, performs the tool’s task, and returns a string. - See other tools for reference. - **Utilities**: `utils/` - The `utils/` directory contains utility functions used across the codebase. - **Additional Details**: {internallinenumbers*} - The agent is very good at automatically utilizing the right available tools at the right time. So do not have an agentic flow that explicitly forces a tool’s usage. {internallinenumbers*} - Common tools, such as file editing and bash commands, are easy for the agent to recognize and use appropriately. However, more complex and niche tools may require explicit instructions in the prompt. {internallinenumbers*} - Tools should be designed to be as general as possible, ensuring they work across any GitHub repository. Avoid hardcoding repository-specific details or behaviors (e.g., paths). - Do not use ’while True’ loops in the agent’s code. This can cause the agent to get stuck and not respond. {internallinenumbers*} - Verify the implementation details of helper functions prior to usage to ensure proper integration and expected behavior. {internallinenumbers*} - Do not install additional packages or dependencies directly. Update `requirements.txt` if new dependencies are required and install them using `pip install -r requirements.txt`. Here is the implementation of the coding agent. # Coding Agent Implementation ----- Coding Agent Implementation Start ----- {code} ----- Coding Agent Implementation End ----- {internallinenumbers*} Your task is to identify ONE detailed plan that would improve the agent’s coding ability. The improvement should not be specific to any particular GitHub issue or repository. Here is the log for the coding agent trying to solve the GitHub issues but failed. # Agent Running Log ----- Agent Running Log Start ----- {md_log} ----- Agent Running Log End ----- # GitHub Issue The GitHub issue that the agent is trying to solve. ----- GitHub Issue Start ----- {github_issue} ----- GitHub Issue End ----- # Predicted Patch The agent’s predicted patch to solve the issue. ----- Predicted Patch Start ----- {predicted_patch} ----- Predicted Patch End ----- # Private Test Patch {internallinenumbers*} SWE-bench’s official private tests to detect whether the issue is solved. This is not available to the agent during evaluation. The agent should try to implement its own tests. ----- Private Test Patch Start ----- {test_patch} ----- Private Test Patch End ----- # Issue Test Results The test results from SWE-bench using the above official private tests. ----- Issue Test Results Start ----- {eval_log} ----- Issue Test Results End ----- Respond precisely in the following format including the JSON start and end markers: ```json <JSON> ``` In <JSON>, provide a JSON response with the following fields: {internallinenumbers*} - "log_summarization": Analyze the above logs and summarize how the agent tried to solve the GitHub issue. Note which tools and how they are used, the agent’s problem-solving approach, and any issues encountered. {internallinenumbers*} - "potential_improvements": Identify potential improvements to the coding agent that could enhance its coding capabilities. Focus on the agent’s general coding abilities (e.g., better or new tools usable across any repository) rather than issue-specific fixes (e.g., tools only usable in one framework). All necessary dependencies and environment setup have already been handled, so do not focus on these aspects. {internallinenumbers*} - "improvement_proposal": Choose ONE high-impact improvement from the identified potential improvements and describe it in detail. This should be a focused and comprehensive plan to enhance the agent’s overall coding ability. {internallinenumbers*} - "implementation_suggestion": Referring to the coding agent’s summary and implementation, think critically about what feature or tool could be added or improved to best implement the proposed improvement. If the proposed feature can be implemented by modifying the existing tools, describe the modifications needed, instead of suggesting a new tool. {internallinenumbers*} - "problem_description": Phrase the improvement proposal and implementation suggestion as a GitHub issue description. It should clearly describe the feature so that a software engineer viewing the issue and the repository can implement it. {internallinenumbers*} Your response will be automatically parsed, so ensure that the string response is precisely in the correct format. Do NOT include the `<JSON>` tag in your output.
@@
Additional Prompt for o3-mini Tool Use and Polyglot ### DOC: tool function schema Carefully consider whether to add/enhance the current tool or edit the workflow in forward() {internallinenumbers*} Pay special attention to making sure that "required" and "type" are always at the correct level of nesting. For example, "required" should be at the same level as "properties", not inside it. Make sure that every property, no matter how short, has a type and description correctly nested inside it. {internallinenumbers*} Other arguments than you have seen are not permitted. For example, in "edit_line_ranges" with "type": "array", arguments like "minItems" and "maxItems" are not permitted. ... Here is the log for the coding agent trying to solve a programming task. {internallinenumbers*} A task is in one programming language, but the coding agent needs to deal with different languages including C++, Go, Java, JavaScript, Python, and Rust.
A.4 Pseudocode
This is the pseudocode of the DGM algorithm, described in Section˜3.

### v2

@@ 6 Conclusion and Limitations
We demonstrate that the DGM can autonomously achieve performance on par with openly available solutions. However, it still falls short of closed-source SoTA SWE-bench solutions. An open question is whether running the DGM for longer would continue to yield performance gains and eventually surpass closed-source solutions. These closed-source solutions often rely on elaborately handcrafted techniques developed by teams of highly skilled experts. Since FMs have yet to match the capabilities of such experts (e.g., in reasoning), the DGM currently requires extensive compute to discover improvements. A single run of the DGM on SWE-bench, as presented in Section˜4, takes about 2 weeks and incurs significant API costs (Section˜E.1). We hypothesize that further progress will require more efficient use of computational resources and the development of better reasoning skills.
@@
Since this version of the DGM is mainly powered by FMs, it is inherently limited by the capabilities of the underlying FM. Hence, an exciting future direction is to extend self-modification beyond just prompts or FM workflows, to include more computationally intensive methods, such as rewriting its own training script to update the FM itself. While this version of the DGM focuses on coding, AI systems are increasingly applied across a wide range of domains (e.g., computer vision, creative writing). Another promising extension is to develop self-improving AI systems capable of enhancing themselves beyond just the coding domain. A key assumption in this work is that coding benchmarks are a good reflection of the agent’s ability to self-improve, since the self-modification task requires the agent to modify its own codebase. However, one could envision an alternative approach that co-evolves the target task distribution (Faldor et al., 2025; Wang et al., 2023c), thereby removing the constraint of self-improvement being tied to a single objective, as in true open-ended processes. Appendix˜I presents additional potential directions for future work. As we continue to explore this powerful technology, we must also keep safety front and center, as discussed in Section˜5.
@@
In conclusion, the DGM represents a significant step toward the automation of AI development through self-improving systems capable of editing their own codebase. While current limitations in compute and reasoning constrain its full potential, continued advances in FMs and infrastructure may unlock more powerful and general-purpose self-improvements. Provided that the safety concerns are carefully navigated (Section˜5), the future of self-improving AI systems and AI-Generating Algorithms (Clune, 2019) holds immense promise to open-endedly evolve AI, continually rewriting or retraining itself in pursuit of greater capabilities aligned with human values.
Acknowledgments
@@
This research was supported by the Vector Institute, the Canada CIFAR AI Chairs program, a grant from Schmidt Futures, an NSERC Discovery Grant, and a generous donation from Rafael Cosman. Resources used in preparing this research were provided, in part, by the Province of Ontario, the Government of Canada through CIFAR, and companies sponsoring the Vector Institute (https://vectorinstitute.ai/partnerships/current-partners/). Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the sponsors. We also thank Aaron Dharna, Ben Norman, Cédric Colas, Sam Devlin, and Shyam Sudhakaran for insightful discussions and feedback.
Appendix
Appendix A Additional Results
A.1 Baselines on SWE-bench
Figure 5: DGM without self-improving agents. Keeping the meta-agent that is modifying and producing the next coding agents the same, DGM w/o self-improve is unable to continuously improve over time. (Left) Archive of coding agents generated during the DGM w/o self-improve run on SWE-bench. Each node represents a coding agent, with node 0 corresponding to the base agent. Node color indicates performance on SWE-bench (percentage of solved tasks), while border color reflects the number of tasks for which the agent was evaluated. Edges show which agents self-modified to produce the offsprings. (Right) Progress plot of the DGM w/o self-improve on SWE-bench. The light green line shows the average score of all agents possessing basic codebase-editing functionality. The green line tracks the best score achieved by any agent in the archive at each iteration. The dark line shows the lineage of the final best-discovered agent and its precursor nodes.
Figure 6: DGM without open-ended exploration. Removing the archive, DGM w/o open-ended exploration always uses the most recent agent to self-modify and makes very little progress on SWE-bench. (Left) Archive of coding agents generated during the DGM w/o open-ended exploration run on SWE-bench. Each node represents a coding agent, with node 0 corresponding to the base agent. Node color indicates performance on SWE-bench (percentage of solved tasks), while border color reflects the number of tasks for which the agent was evaluated. Edges show which agents self-modified to produce the offsprings. (Right) Progress plot of the DGM w/o open-ended on SWE-bench. The orange line shows the average score of all agents possessing basic codebase-editing functionality. The light orange line tracks the best score achieved by any agent in the archive at each iteration. The dark line shows the lineage of the final best-discovered agent and its precursor nodes.
A.2 Generality across models on Polyglot
Figure 7: Transfer between Models on Polyglot
In addition to testing the transfer models on SWE-bench (see Section˜4.4, Figure˜2), we also present the transfer results on Polyglot in this section. On Polyglot (Figure˜7), where the DGM was run with o3-mini, we replaced the FM with Claude 3.5 Sonnet (New) or Claude 3.7 Sonnet, and evaluated on the full benchmark (Figure˜4, Middle). With Claude 3.5 Sonnet (New), the initial agent achieved 32.0% and the DGM-discovered agent 33.3%. With Claude 3.7 Sonnet, the initial agent achieved 35.6% and the DGM-discovered agent 36.8%. These results suggest that the DGM yields improvements that generalize across FMs, rather than being tightly coupled to the specific FM used during its run (Figure˜4).
A.3 Ablation of Parent Selection
To further study the impact of the parent selection mechanism in DGM, we introduce DGM Greedy. DGM Greedy always selects the best-performing node as the parent to branch from, rather than giving every node a non-zero probability of being branched off (roughly proportional to their performance score and number of children) as in this implementation of the DGM (Section˜C.2). This ablation replicates the approach of Robeyns et al. (2025) in this setting. As shown in Table˜1, DGM Greedy achieves 39.7% and 30.0% on SWE-bench and Polyglot, respectively, compared to 50.0% and 38.0% by this implementation of DGM. These results demonstrate that allowing all solutions in the archive to serve as potential stepping stones can lead to greater improvements over time, underscoring the importance of open-ended exploration.
Table 1: Comparison of DGM, its ablations, and baselines on SWE-bench and Polyglot benchmarks.
Method SWE-bench Polyglot DGM 50.0% 38.0% DGM w/o Open-ended exploration 23.0% 14.0% DGM w/o Self-improve 39.0% 28.0% DGM Greedy 39.7% 30.0%
A.4 Additional Statistics of DGM runs
Percentage of Generated Agents with Basic Code-Editing Functionality. To gain deeper insights into the DGM process, we analyze the percentage of generated agents that possess basic code-editing functionality on the SWE-bench benchmark. As shown in Table˜2, DGM exhibits the highest percentage of producing agents with basic codebase-editing functionality. These results highlight the effectiveness of both open-ended exploration and self-improvement components in the DGM, where open-ended exploration enables the search to escape local optima, while self-improvement enhances the ability to generate better agents.
Table 2: Percentage of generated agents with basic code-editing functionality on SWE-bench.
Method Percentage with Basic Code-Editing Functionality DGM 51.3% DGM w/o Open-ended exploration 32.5% DGM w/o Self-improve 32.5%
Stability of DGM Runs. To evaluate the stability of DGM, we run the DGM algorithm three times on the Polyglot benchmark and analyze the variance in performance. The DGM achieved a mean accuracy of 40.7% with a standard deviation of 2.3%, indicating that the DGM can achieve consistent and reproducible results across runs.
Appendix B Additional Related Work
Open-Endedness (part 2). Early approaches to open-endedness explored different mechanisms to balance learnability and interestingness. Quality-diversity algorithms sought to illuminate vast solution spaces with diverse, high-performing behaviors (Pugh et al., 2016; Chatzilygeroudis et al., 2021; Mouret & Clune, 2015; Nguyen et al., 2015). Other methods emphasized goal-directed exploration (Ecoffet et al., 2019; 2021; Schaul et al., 2015; Andrychowicz et al., 2017; Eysenbach et al., 2018), intrinsic motivation (Lehman & Stanley, 2011; Oudeyer et al., 2007; Li et al., 2014; Pathak et al., 2017), or learning progress frameworks (Kanitscheider et al., 2021; Gaven et al., 2025; Baranes & Oudeyer, 2013; Colas et al., 2019; 2022b; Jiang et al., 2021; Dennis et al., 2020; Schmidhuber, 2008; 2013; Kompella et al., 2017). More recently, large-scale foundation models (FMs) (Brown et al., 2020; Radford et al., 2019) have emerged as powerful proxies for human notions of interestingness (Zhang et al., 2024b; Faldor et al., 2025; Sancaktar et al., 2025) and effective mutation operators to propose novel solutions in code (Romera-Paredes et al., 2024; Novikov et al., 2025; Lehman et al., 2023; Faldor et al., 2025; Hu et al., 2025). FMs can guide autotelic agents (Colas et al., 2022b; 2023; a), model human preferences for quality and diversity (Bradley et al., 2024; Ding et al., 2024; Wang et al., 2023b; Klissarov et al., 2023; 2024; Samvelyan et al., 2024; Lim et al., 2024; Havrilla et al., 2024a), design reward functions (Wiering & Schmidhuber, 1997; Wang et al., 2023a; Ma et al., 2023; Faldor et al., 2025), create simulated environments (Sudhakaran et al., 2023; Nasir & Togelius, 2023; Aki et al., 2024; Nasir et al., 2024; Bruce et al., 2024; Parker-Holder et al., 2024; Schmidhuber, 2013), drive ever-evolving multi-agent dynamics (Dharna et al., 2024; Zhou et al., 2025), search diverse ambulating robot morphologies (Lehman et al., 2023), and search expansive solution spaces for benchmark or objective optimization (Lange et al., 2024; Zhang et al., 2024b; Faldor et al., 2025; Hu et al., 2025; Lu et al., 2024b; Romera-Paredes et al., 2024; Fernando et al., 2024; Lu et al., 2024a; Khan et al., 2024; Lu et al., 2025; Liu et al., 2024; Novikov et al., 2025).
Program Synthesis. Program synthesis (Alur et al., 2018; Polozov & Gulwani, 2015; Buchi & Landweber, 1990; Gulwani, 2011; Ellis et al., 2021) seeks to generate code meeting external specifications such as input-output examples or logical formulas. Hybrid approaches combine symbolic methods with neural or FM guidance: for instance, Li et al. (2024) uses LLM suggestions to steer symbolic search in SyGuS settings, improving over pure enumeration. Barke et al. (2024) blends LLM completions with a learned surrogate model to guide synthesis in DSLs. Shi et al. (2023) uses neural policies to build higher-order and lambda abstractions during search, outperforming both LLM-only and symbolic baselines on list manipulation tasks. The DGM differs in focusing not just on producing programs for external tasks, but also on agent self-modification, rewriting its own implementation to improve its capacity for future self-improvement.
Inspiration from Darwinian Evolution. This work is heavily inspired by the mechanisms of Darwinian evolution (Darwin, 2023), notably variation (mutation), selection, and the preservation of lineages (stepping stones), and brings them into the realm of self-modifying coding agents. In DGM, an archive of past agent versions is maintained, from which parent agents are sampled; then mutations (i.e., code edits) generate new offspring agents, which are empirically evaluated on coding benchmarks; successful ones are added to the archive, enabling parallel exploration of multiple evolutionary trajectories (Section˜3). This mirrors how biological evolution (Edwards, 2000; Wright, 1932) preserves genetic diversity (Mayr, 1982), leverages variation (Kimura, 1979), and uses natural selection to retain beneficial mutations (Dobzhansky, 1970).
Appendix C Algorithmic Details
C.1 Initial Coding Agent
@@
C.2 Parent Selection
At each DGM iteration, we select a subset of agents from the archive as parents to self-modify and produce new child agents (Section˜3). The details of the parent selection process, inspired by Ecoffet et al. (2019), are outlined below. Future work could explore alternative search and exploration methods (Coulom, 2006; Silver et al., 2016; Herr et al., 2025).
@@ C.2 Parent Selection
C.3 Self-Improve Prompts
In every iteration, an agent is chosen to self-modify. For this step, the current implementation of DGM leverages a separate FM to diagnose improvement opportunities, which are then passed to the coding agent for implementation. To identify the next potential improvement, the evaluation logs from the chosen agent’s performance on the coding benchmark are fed into the FM. While the current implementation relies on an additional diagnostic step, there are no fundamental limitations preventing the DGM from autonomously analyzing its own performance, generating improvement hypotheses, and implementing subsequent modifications. We chose to use OpenAI’s o1 model (Team et al., 2024) for this purpose, as we found it had better reasoning capabilities than other FMs available at the time of our experiments. The prompts used for this diagnosis are provided below. From the diagnostic output, the implementation_suggestion and problem_description are extracted and passed to the coding agent as its self-modification task.
@@ C.3 Self-Improve Prompts
Self-Improve Prompt # Coding Agent Summary - **Main File**: `coding_agent.py` - Primary Class: `AgenticSystem` - The `forward()` function is the central entry point. {internallinenumbers*} - Prompts are located either within the `forward()` function or in the `prompts/` directory. - **Tools**: `tools/` {internallinenumbers*} - The `tools/` directory contains various tools that LLMs can use to perform specific tasks. {internallinenumbers*} - Each tool must have a `tool_info()` function that returns a JSON object containing ’name’, ’description’, and ’input_schema’. The ’input_schema’ should be a JSON object containing ’type’, ’properties’, and ’required’. {internallinenumbers*} - Each tool must have a `tool_function()` function that takes the arguments defined in input_schema, performs the tool’s task, and returns a string. - See other tools for reference. - **Utilities**: `utils/` {internallinenumbers*} - The `utils/` directory contains utility functions used across the codebase. - **Additional Details**: {internallinenumbers*} - The agent is very good at automatically utilizing the right available tools at the right time. So do not have an agentic flow that explicitly forces a tool’s usage. {internallinenumbers*} - Common tools, such as file editing and bash commands, are easy for the agent to recognize and use appropriately. However, more complex and niche tools may require explicit instructions in the prompt. {internallinenumbers*} - Tools should be designed to be as general as possible, ensuring they work across any GitHub repository. Avoid hardcoding repository-specific details or behaviors (e.g., paths). {internallinenumbers*} - Do not use ’while True’ loops in the agent’s code. This can cause the agent to get stuck and not respond. {internallinenumbers*} - Verify the implementation details of helper functions prior to usage to ensure proper integration and expected behavior. {internallinenumbers*} - Do not install additional packages or dependencies directly. Update `requirements.txt` if new dependencies are required and install them using `pip install -r requirements.txt`. Here is the implementation of the coding agent. # Coding Agent Implementation ----- Coding Agent Implementation Start ----- {code} ----- Coding Agent Implementation End ----- {internallinenumbers*} Your task is to identify ONE detailed plan that would improve the agent’s coding ability. The improvement should not be specific to any particular GitHub issue or repository. {internallinenumbers*} Here is the log for the coding agent trying to solve the GitHub issues but failed. # Agent Running Log ----- Agent Running Log Start ----- {md_log} ----- Agent Running Log End ----- # GitHub Issue The GitHub issue that the agent is trying to solve. ----- GitHub Issue Start ----- {github_issue} ----- GitHub Issue End ----- # Predicted Patch The agent’s predicted patch to solve the issue. ----- Predicted Patch Start ----- {predicted_patch} ----- Predicted Patch End ----- # Private Test Patch {internallinenumbers*} SWE-bench’s official private tests to detect whether the issue is solved. This is not available to the agent during evaluation. The agent should try to implement its own tests. ----- Private Test Patch Start ----- {test_patch} ----- Private Test Patch End ----- # Issue Test Results The test results from SWE-bench using the above official private tests. ----- Issue Test Results Start ----- {eval_log} ----- Issue Test Results End ----- {internallinenumbers*} Respond precisely in the following format including the JSON start and end markers: ```json <JSON> ``` In <JSON>, provide a JSON response with the following fields: {internallinenumbers*} - "log_summarization": Analyze the above logs and summarize how the agent tried to solve the GitHub issue. Note which tools and how they are used, the agent’s problem-solving approach, and any issues encountered. {internallinenumbers*} - "potential_improvements": Identify potential improvements to the coding agent that could enhance its coding capabilities. Focus on the agent’s general coding abilities (e.g., better or new tools usable across any repository) rather than issue-specific fixes (e.g., tools only usable in one framework). All necessary dependencies and environment setup have already been handled, so do not focus on these aspects. {internallinenumbers*} - "improvement_proposal": Choose ONE high-impact improvement from the identified potential improvements and describe it in detail. This should be a focused and comprehensive plan to enhance the agent’s overall coding ability. {internallinenumbers*} - "implementation_suggestion": Referring to the coding agent’s summary and implementation, think critically about what feature or tool could be added or improved to best implement the proposed improvement. If the proposed feature can be implemented by modifying the existing tools, describe the modifications needed, instead of suggesting a new tool. {internallinenumbers*} - "problem_description": Phrase the improvement proposal and implementation suggestion as a GitHub issue description. It should clearly describe the feature so that a software engineer viewing the issue and the repository can implement it. {internallinenumbers*} Your response will be automatically parsed, so ensure that the string response is precisely in the correct format. Do NOT include the `<JSON>` tag in your output.
While the private test patch is included in the diagnosis prompt, one could imagine that an easy way for the agent to score well on the task would be to hard code the solutions. However, from manual inspection of the logs and agent modifications, we have not observed any problematic logic or behavior indicative of memorization or overfitting to specific private test cases. Furthermore, improvements discovered on one task or LLM model often generalize well to others (Figure˜4), showing that these improvements are not task-specific but instead reflect general changes that enhance overall performance.
@@
Additional Prompt for o3-mini Tool Use and Polyglot ### DOC: tool function schema {internallinenumbers*} Carefully consider whether to add/enhance the current tool or edit the workflow in forward() {internallinenumbers*} Pay special attention to making sure that "required" and "type" are always at the correct level of nesting. For example, "required" should be at the same level as "properties", not inside it. {internallinenumbers*} Make sure that every property, no matter how short, has a type and description correctly nested inside it. {internallinenumbers*} Other arguments than you have seen are not permitted. For example, in "edit_line_ranges" with "type": "array", arguments like "minItems" and "maxItems" are not permitted. ... Here is the log for the coding agent trying to solve a programming task. {internallinenumbers*} A task is in one programming language, but the coding agent needs to deal with different languages including C++, Go, Java, JavaScript, Python, and Rust.
C.4 Pseudocode
This is the pseudocode of the DGM algorithm, described in Section˜3: